Ela is twice as old as her friend kath. kath is 7 years older than raquel. Ten years ago, ela was three times as old as raquel. How old is Kath now?
Answer by lwsshak3(11628)

Ela is twice as old as her friend kath. kath is 7 years older than raquel. Ten years ago, ela was three times as old as raquel. How old is Kath now?
***
now:
x=requel's age
x+7=kath's age
2(x+7)=ela's age
..
10 yrs ago:
x-10=requel's age
x+7-10=x-3=kath's age
2(x+7)-10=2x+4=ela's age
2x+4=3(x-10)
2x+4=3x-30
x=34
x+7=41
How old is Kath now? 47